# Generative Large-Scale Pre-trained Models for Automated Ad Bidding Optimization

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: conference: Preprint; April 4; 2025isbn: ccs: Information systems Computational advertising
@@ 1. Introduction
Online advertising has become a vital revenue engine for internet platforms, spurring the development of intelligent ad delivery systems. As the online advertising market continues to expand, advertisers are raising their expectations for both the effectiveness and efficiency of their campaigns (Evans, 2009). A core competency of advertising platforms lies in how intelligently their auto-bidding systems help advertisers maximize marketing value under constraints such as budgets and performance objectives (Yang et al., 2019; Wen et al., 2022; Wang et al., 2023). In practical industrial scenarios, advertisers exhibit highly heterogeneous and dynamic preferences (Barford et al., 2014) – some prioritize brand exposure while others focus on conversion rates or cost efficiency (Lee et al., 2012). Consequently, auto-bidding systems must optimize general performance (Aggarwal et al., 2024; Deng et al., 2021) while accommodating flexible support for a wide range of objectives and constraints (Ou et al., 2023), adapting swiftly to shifting business demands and market environments.
Recent advances in reinforcement learning (Jin et al., 2018; Ye et al., 2019; Mou et al., 2022) and generative modeling (Li et al., 2024b) have significantly transformed the landscape of bidding algorithms, which empower the development of more intelligent (Bijalwan et al., 2025), adaptive, and personalized solutions tailored to complex and varied industrial needs of modern advertisers (Jiang, 2025). In particular, conditional generative approaches, leveraging robust architectures such as transformers (Li et al., 2024b; Jiang et al., 2025; Lei et al., 2024) and diffusion models (Guo et al., 2024), enable direct generation of bidding actions or entire bidding trajectories by encoding advertiser preferences as vectorized conditions. This paradigm bypasses the limitations of traditional Markov Decision Process (MDP) frameworks and allows for more flexible real-time decision-making. Inspired by the success of large-scale foundation models like GPT-4 (Achiam et al., 2023), DeepSeek (Liu et al., 2024), and Stable Diffusion (Rombach et al., 2022), the auto-bidding field is now moving towards foundation model-based methods, where universal decision policies are learned from massive datasets, enabling scalable and adaptive bidding systems capable of addressing a broad spectrum of advertiser needs.
@@
To address these challenges fundamentally, we propose GRAD (Generative Reward-driven AD-bidding), a novel and scalable generative auto-bidding architecture. Built upon the principle of the scaling law—where model performance scales exponentially with computational resources—GRAD is designed to achieve both adaptability and deployment reliability in industrial auto-bidding systems. This architecture provides a robust solution to offline-online discrepancy and suboptimal action space limitations through synergistic innovations. Specifically, the Action-MoE Module employs a Mixture-of-Experts (MoE) mechanism (Huang et al., 2025) for constrained exploration in high-dimensional action spaces, dynamically generating novel strategies beyond historically observed actions while inherently mitigating feature distribution shifts through domain-informed constraint simulation. Complementing this, the Causal Transformer Value Estimator performs counterfactual inference to evaluate the reward of unexecuted actions under complex advertiser constraints, enabling precise and objective-aligned optimization. Together, this decoupled exploration-optimization paradigm—powered by scalable MoE and deep causal reasoning—has been successfully deployed across Meituan’s advertising platform, demonstrating a sustainable pathway for reliable generative auto-bidding in real-world industrial environments.
@@ 2.1. Auto-Bidding Problem
The optimization problem above involves a high-dimensional decision space and uncertain auction outcomes. Previous work (He et al., 2021; Su et al., 2024) reformulates it as a linear program and derives closed-form bidding rules. In the presence of a single CPC constraint, the optimal bidding $b^{*}_{i}$ can be calculated by:
@@ 2.2. Auto-bidding Decision Process
To address this complexity, we adopt the Transformer paradigm (Chen et al., 2021; Roberts, 2024; Hu et al., 2023) to model sequential dependencies in the bidding process, which we formalize as a MDP. The MDP is specified by the tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$ where $\mathcal{S}$, $\mathcal{A}$, $\mathcal{T}$, and $\mathcal{R}$ denote the state space, action space, state transition function, and reward function, respectively, and $\gamma$ is the discount factor. At each decision cycle $t$, the agent observes the current state $s_{t}\in\mathcal{S}$, selects an action $a_{t}\in\mathcal{A}$, transitions to the next state $s_{t+1}=\mathcal{T}(s_{t},a_{t})$ according to the environment dynamics, and receives an immediate reward $r_{t}$ that quantifies the value generated by the action.
@@
  • Reward ($r_{t}$): The immediate value accrued from impressions obtained between decision cycles $t-1$ and $t$. This reward is typically measured as the sum of the realized value (e.g., clicks, conversions, or revenue) generated by the set of impressions acquired in that interval, providing direct feedback to guide policy optimization.
@@ 3. Method
Figure 1. Overall architecture: (a) a Causal Transformer module based on the decoder-only architecture of Transformer(see §3.1); (b) a Value Estimator (see §3.2); (c) an Action MoE module, as described in §3.3.
The proposed framework, GRAD, establishes a novel foundation for adaptive auto-bidding, which is visually detailed in Figure 1. GRAD employs a Causal Transformer (CT) (Chen et al., 2021) to model complex temporal dependencies in state-action histories. To address critical challenges in online advertising, it integrates two specialized modules: ActionMoE , which enhances exploration efficacy and policy robustness through a learnable MoE mechanism; and the Value Estimator, which incorporates explicit constraints (e.g., budget pacing and CPC targets) directly into the reward signal. This synergistic combination enables the GRAD framework to effectively balance exploration-exploitation trade-offs, dynamically satisfy complex operational constraints, and generate high-quality bidding decisions in real-time ad bidding.
@@ 3.1. Causal Transformer
where $\oplus$ denotes vector concatenation, $E_{g}$, $E_{s}$, and $E_{a}$ are linear embedding layers for the RTG, state, and previous action, respectively, $\text{PE}(t)$ (Kazemnejad et al., 2023) denotes the positional encoding at timestep $t$.
@@ 3.3. ActionMoE
To tackle the challenge of action space design under hard constraints, particularly in complex or high-dimensional environments where conventional exploration may be inadequate or unsafe, we propose ActionMoE—a lightweight exploration module for transformer-based policies. Drawing on DeepSeekMoE (Dai et al., 2024) and Switch Transformers (Fedus et al., 2022), ActionMoE employs fine-grained expert segmentation and shared expert isolation. Through a mixture-of-experts and adaptive gating, it enhances exploration, increases action diversity, and maintains stable, safe training (Song et al., 2024; Mu and Lin, 2025; Dey and Henze, 2024). As depicted in Figure 1, the ActionMoE can be organisazed as follows:
@@
11:  $\quad\mathbf{a}_{t}^{\ast}\leftarrow\omega_{m}\mathbf{a}^{m}_{t-1}+\mathbf{U}_{t}$triangleright $\omega_{m}$: learnable weights
@@
where Auxiliary Loss (AUX) (Wang et al., 2024) penalizes imbalanced expert usage and $\mathbf{p}$ represents the routing probabilities assigned to each expert, and $\mathbf{u}$ denotes the observed utilization frequency of experts within a mini-batch. The anchoring term $\|\mathbf{h}_{t}-\mathbf{h}_{t}^{\text{shared}}\|_{2}^{2}$ aligns the hidden state with the shared expert output $\mathbf{h}_{t}^{\text{shared}}=\operatorname{FFN}_{\text{shared}}(\mathbf{h}_{t})$. The set $\lambda_{\text{aux}}$ represents hyperparameters that balance the relative importance of primary and auxiliary loss components.
@@ 4.1. Setup
Datasets. Unlike previous work that relies on private bidding logs from closed-source systems, we adopt AuctionNet, a publicly available large-scale bidding dataset released by Alibaba (Su et al., 2024). To our knowledge, it is the largest public dataset in this domain, containing over 500 million auction records. A sparse variant is also provided to simulate more challenging scenarios. The key statistics are summarized in Table 4 from the Appendix A.1.
Evaluation Metrics. Following GAVE (Gao et al., 2025), we evaluate performance using a composite score that balances total value and constraint satisfaction. Specifically, we define:
@@ 4.2. Comparative Evaluation
  • DiffBid (Guo et al., 2024): Models bidding sequences using diffusion processes for trajectory generation.
  • USCB (He et al., 2021): Online RL-based approach that dynamically adjusts bidding parameters via a unified optimization framework.
  • CQL (Kumar et al., 2020): Conservative Q-learning algorithm that penalizes overestimation in offline settings.
  • IQL (Kostrikov et al., [n. d.]): Utilizes expectile regression to improve policies without relying on value extrapolation.
  • BCQ (Fujimoto et al., 2019): Restricts action selection via a learned generative model to remain close to the dataset policy.
  • DT (Chen et al., 2021): Uses a transformer architecture to model decision trajectories via behavior cloning.
  • CDT (Liu et al., 2023): Offline RL approach for learning safe policies under explicit constraint satisfaction.
  • GAS (Li et al., 2024b): Enhances DT with post-training Monte Carlo Tree Search (MCTS) for policy refinement.
  • GAVE (Gao et al., 2025): Introduces constrained exploration and RTG-based budget control within a Causal Transformer framework.
Comparison Design. All methods are evaluated on the AuctionNet dataset and its sparse variant under varying budget constraints (50%, 75%, 100%, 125%, 150% of the maximum). Performance is measured using the composite score metric described in Section 4.1, which balances total value and constraint satisfaction. During training, we utilize the PyTorch on two NVIDIA A100 GPUs. For more detailed hyperparameters, please refer to Table  5 from Appendix A.2.
@@ 5.2. Deployment
[…]
@@ 5.3. Online A/B Test Results
[…]
@@ 6.1. Offline Reinforcement Learning
Offline reinforcement learning (RL) learns optimal policies from static datasets (Schweighofer et al., 2022; Lepak and Wawrzyński, 2024), eliminating costly/dangerous online interaction. Unlike traditional RL requiring active exploration, it extracts behaviors solely from pre-collected data (Ball et al., 2023; Levine et al., 2020), proving invaluable in exploration-constrained domains.
Recent advances in this field have introduced innovative techniques to address key challenges such as distributional shift and overestimation bias (Aggarwal et al., 2025). Conservative Q-learning (CQL) (Kumar et al., 2020) mitigates value overestimation by penalizing out-of-distribution actions, while Batch-Constrained Q-learning (BCQ) (Fujimoto et al., 2019) restricts policy updates to actions supported by the dataset. Implicit Q-learning (IQL)  (Kostrikov et al., [n. d.]) further improves stability by learning value functions without explicit policy constraints. Additionally, Decision Transformer (Chen et al., 2021) leverages sequence modeling to capture long-range dependencies in offline data, and IDQL (Hansen-Estruch et al., 2023) integrates diffusion models with implicit Q-learning to enhance policy robustness. These approaches collectively advance the scalability and reliability of offline RL in complex, real-world settings.
@@ 6.2. Auto-bidding
Auto-bidding systems have become a cornerstone of programmatic advertising (He et al., 2024), enabling automated real-time bidding (Sayedi, 2018; Liu et al., 2022; Cai et al., 2025) (RTB) for ad placements while optimizing key performance indicators (KPIs) such as click-through rates (CTR), conversions, or return on ad spend (ROAS) under strict budget constraints. These systems leverage machine learning and optimization techniques to adjust bids in complex, competitive auction environments dynamically.
Reinforcement learning (RL) has gained significant traction in optimizing auto-bidding strategies (Korenkevych et al., 2024; Li et al., 2024a; Chen et al., 2023). The field has evolved from early simulated bidding environments to (He et al., 2021)’s unified approach handling multiple constraints, followed by (Jin et al., 2018)’s multi-agent competitive settings and (Mou et al., 2022)’s performance improvements over rule-based strategies. Most recently, (Wen et al., 2022) deployed a cooperative multi-agent framework incorporating richer market information. Despite demonstrated successes, these RL approaches continue to face challenges, including sample inefficiency, training instability, and adaptation difficulties in dynamic auction environments.
@@ 6.3. Generative Models
Generative models aim to capture the underlying data distribution of a given dataset or model conditional probability distributions between variables. Recently, transformer-based architectures have been adapted to model complex DMP through autoregressive generative frameworks, as demonstrated in models like GAVE (Gao et al., 2025) and GAS (Li et al., 2024b). Concurrently, diffusion models have emerged as a powerful paradigm for conditional sample generation. These models iteratively refine samples via a reverse denoising process conditioned on inputs, with representative examples including DiffBid (Guo et al., 2024).

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025conference: KDD 2026; August; 2026ccs: Information systems Computational advertising
@@ 1. Introduction
Online advertising has become a vital revenue engine for internet platforms, spurring the development of intelligent ad delivery systems. As the online advertising market continues to expand, advertisers are raising their expectations for both the effectiveness and efficiency of their campaigns (evans2009online). A core competency of advertising platforms lies in how intelligently their auto-bidding systems help advertisers maximize marketing value under constraints such as budgets and performance objectives (yang2019aiads; wen2022cooperative; wang2023adversarial; liu2024bird). In practical industrial scenarios, advertisers exhibit highly heterogeneous and dynamic preferences (barford2014adscape; liu2025sigma) – some prioritize brand exposure while others focus on conversion rates or cost efficiency (lee2012estimating). Consequently, auto-bidding systems must optimize general performance (aggarwal2024auto; deng2021towards) while accommodating flexible support for a wide range of objectives and constraints (ou2023survey), adapting swiftly to shifting business demands and market environments.
Recent advances in reinforcement learning (jin2018real; ye2019deep; mou2022sustainable) and generative modeling (li2024gas; liu2025deepfake) have significantly transformed the landscape of bidding algorithms, which empower the development of more intelligent (bijalwan2025unveiling), adaptive, and personalized solutions tailored to complex and varied industrial needs of modern advertisers (jiang2025automated). In particular, conditional generative approaches, leveraging robust architectures such as transformers (li2024gas; jiang2025optimal; lei2024generative; zhang2025diffusion) and diffusion models (guo2024generative; zhang2506timemaster), enable direct generation of bidding actions or entire bidding trajectories by encoding advertiser preferences as vectorized conditions. This paradigm bypasses the limitations of traditional Markov Decision Process (MDP) frameworks and allows for more flexible real-time decision-making. Inspired by the success of large-scale foundation models like GPT-4 (achiam2023gpt), DeepSeek (liu2024deepseek), and Stable Diffusion (rombach2022high), the auto-bidding field is now moving towards foundation model-based methods, where universal decision policies are learned from massive datasets, enabling scalable and adaptive bidding systems capable of addressing a broad spectrum of advertiser needs.
@@
To address these challenges fundamentally, we propose GRAD (Generative Reward-driven AD-bidding), a novel and scalable generative auto-bidding architecture. Built upon the principle of the scaling law—where model performance scales exponentially with computational resources—GRAD is designed to achieve both adaptability and deployment reliability in industrial auto-bidding systems. This architecture provides a robust solution to offline-online discrepancy and suboptimal action space limitations through synergistic innovations. Specifically, the Action-MoE Module employs a Mixture-of-Experts (MoE) mechanism (huang2025moe) for constrained exploration in high-dimensional action spaces, dynamically generating novel strategies beyond historically observed actions while inherently mitigating feature distribution shifts through domain-informed constraint simulation. Complementing this, the Causal Transformer Value Estimator performs counterfactual inference to evaluate the reward of unexecuted actions under complex advertiser constraints, enabling precise and objective-aligned optimization. Together, this decoupled exploration-optimization paradigm—powered by scalable MoE and deep causal reasoning—has been successfully deployed across Meituan’s advertising platform, demonstrating a sustainable pathway for reliable generative auto-bidding in real-world industrial environments.
@@ 2.1. Auto-Bidding Problem
The optimization problem above involves a high-dimensional decision space and uncertain auction outcomes. Previous work (he2021unified; su2024auctionnet) reformulates it as a linear program and derives closed-form bidding rules. In the presence of a single CPC constraint, the optimal bidding $b^{*}_{i}$ can be calculated by:
@@ 2.2. Auto-bidding Decision Process
To address this complexity, we adopt the Transformer paradigm (chen2021decision; roberts2024powerful; hu2023graph) to model sequential dependencies in the bidding process, which we formalize as a MDP. The MDP is specified by the tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$ where $\mathcal{S}$, $\mathcal{A}$, $\mathcal{T}$, and $\mathcal{R}$ denote the state space, action space, state transition function, and reward function, respectively, and $\gamma$ is the discount factor. At each decision cycle $t$, the agent observes the current state $s_{t}\in\mathcal{S}$, selects an action $a_{t}\in\mathcal{A}$, transitions to the next state $s_{t+1}=\mathcal{T}(s_{t},a_{t})$ according to the environment dynamics, and receives an immediate reward $r_{t}$ that quantifies the value generated by the action.
@@
  • Reward ($r_{t}$): The immediate value accrued from impressions obtained between decision cycles t1t-1 and $t$. This reward is typically measured as the sum of the realized value (e.g., clicks, conversions, or revenue) generated by the set of impressions acquired in that interval, providing direct feedback to guide policy optimization.
@@ 3. Method
Figure 1. Overall architecture: (a) a Causal Transformer module based on the decoder-only architecture of Transformer(see §3.1); (b) a Value Estimator (see §3.2); (c) an Action MoE module, as described in §$3.$
The proposed framework, GRAD, establishes a novel foundation for adaptive auto-bidding, which is visually detailed in Figure 1. GRAD employs a Causal Transformer (CT) (chen2021decision) to model complex temporal dependencies in state-action histories. To address critical challenges in online advertising, it integrates two specialized modules: ActionMoE , which enhances exploration efficacy and policy robustness through a learnable MoE mechanism; and the Value Estimator, which incorporates explicit constraints (e.g., budget pacing and CPC targets) directly into the reward signal. This synergistic combination enables the GRAD framework to effectively balance exploration-exploitation trade-offs, dynamically satisfy complex operational constraints, and generate high-quality bidding decisions in real-time ad bidding.
@@ 3.1. Causal Transformer
where $\oplus$ denotes vector concatenation, $E_{g}$, $E_{s}$, and $E_{a}$ are linear embedding layers for the RTG, state, and previous action, respectively, $\text{PE}(t)$ (kazemnejad2023impact) denotes the positional encoding at timestep $t$.
@@ 3.3. ActionMoE
To tackle the challenge of action space design under hard constraints, particularly in complex or high-dimensional environments where conventional exploration may be inadequate or unsafe, we propose ActionMoE—a lightweight exploration module for transformer-based policies. Drawing on DeepSeekMoE (dai2024deepseekmoe) and Switch Transformers (fedus2022switch), ActionMoE employs fine-grained expert segmentation and shared expert isolation. Through a mixture-of-experts and adaptive gating, it enhances exploration, increases action diversity, and maintains stable, safe training (song2024germ; mu2025comprehensive; dey2024reinforcement). As depicted in Figure 1, the ActionMoE can be organisazed as follows:
@@
11:  $\quad\mathbf{a}_{t}^{\ast}\leftarrow\omega_{m}\mathbf{a}^{m}_{t-1}+\mathbf{U}_{t}$ $\triangleright$ $\omega_{m}$: learnable weights
@@
where Auxiliary Loss (AUX) (wang2024auxiliary) penalizes imbalanced expert usage and $\mathbf{p}$ represents the routing probabilities assigned to each expert, and $\mathbf{u}$ denotes the observed utilization frequency of experts within a mini-batch. The anchoring term $\|\mathbf{h}_{t}-\mathbf{h}_{t}^{\text{shared}}\|_{2}^{2}$ aligns the hidden state with the shared expert output $\mathbf{h}_{t}^{\text{shared}}=\operatorname{FFN}_{\text{shared}}(\mathbf{h}_{t})$. The set $\lambda_{\text{aux}}$ represents hyperparameters that balance the relative importance of primary and auxiliary loss components.
@@ 4.1. Setup
Datasets. Unlike previous work that relies on private bidding logs from closed-source systems, we adopt AuctionNet, a publicly available large-scale bidding dataset released by Alibaba (su2024auctionnet). To our knowledge, it is the largest public dataset in this domain, containing over 500 million auction records. A sparse variant is also provided to simulate more challenging scenarios. The key statistics are summarized in Table 4 from the Appendix A.
Evaluation Metrics. Following GAVE (gao2025generative), we evaluate performance using a composite score that balances total value and constraint satisfaction. Specifically, we define:
@@ 4.2. Comparative Evaluation
  • DiffBid (guo2024generative): Models bidding sequences using diffusion processes for trajectory generation.
  • USCB (he2021unified): Online RL-based approach that dynamically adjusts bidding parameters via a unified optimization framework.
  • CQL (kumar2020conservative): Conservative Q-learning algorithm that penalizes overestimation in offline settings.
  • IQL (kostrikovoffline): Utilizes expectile regression to improve policies without relying on value extrapolation.
  • BCQ (fujimoto2019off): Restricts action selection via a learned generative model to remain close to the dataset policy.
  • DT (chen2021decision): Uses a transformer architecture to model decision trajectories via behavior cloning.
  • CDT (liu2023constrained): Offline RL approach for learning safe policies under explicit constraint satisfaction.
  • GAS (li2024gas): Enhances DT with post-training Monte Carlo Tree Search (MCTS) for policy refinement.
  • GAVE (gao2025generative): Introduces constrained exploration and RTG-based budget control within a Causal Transformer framework.
Comparison Design. All methods are evaluated on the AuctionNet dataset and its sparse variant under varying budget constraints (50%, 75%, 100%, 125%, 150% of the maximum). Performance is measured using the composite score metric described in Section 4.1, which balances total value and constraint satisfaction. During training, we utilize the PyTorch on two NVIDIA A100 GPUs. For more detailed hyperparameters, please refer to Table  5 from Appendix B.
@@ 5.2. Deployment
  • Training. Historical users’ behavior data are aggregated into groups and serve as inputs. The model optimizes bidding parameters under budget and KPI constraints (e.g., CPC). The architecture is based on a Causal Transformer with two key components: (1) ActionMoE, which integrates a shared expert for general feature extraction and domain-specific routed experts with top-$K=1$ routing, and (2) Value Estimator, which enforces joint budget-KPI constraint adherence.
@@ 5.3. Online A/B Test Results
(27) $\text{CPC\_CR}=\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}\left(\overline{\text{CPC}}_{t}\leq\gamma C_{\text{target}}\right)\times 100\%$
@@ 6.1. Offline Reinforcement Learning
Offline reinforcement learning (RL) learns optimal policies from static datasets (schweighofer2022dataset; lepak2024reinforcement), eliminating costly/dangerous online interaction. Unlike traditional RL requiring active exploration, it extracts behaviors solely from pre-collected data (ball2023efficient; levine2020offline), proving invaluable in exploration-constrained domains.
Recent advances in this field have introduced innovative techniques to address key challenges such as distributional shift and overestimation bias (aggarwal2025multi). Conservative Q-learning (CQL) (kumar2020conservative) mitigates value overestimation by penalizing out-of-distribution actions, while Batch-Constrained Q-learning (BCQ) (fujimoto2019off) restricts policy updates to actions supported by the dataset. Implicit Q-learning (IQL)  (kostrikovoffline) further improves stability by learning value functions without explicit policy constraints. Additionally, Decision Transformer (chen2021decision) leverages sequence modeling to capture long-range dependencies in offline data, and IDQL (hansen2023idql) integrates diffusion models with implicit Q-learning to enhance policy robustness. These approaches collectively advance the scalability and reliability of offline RL in complex, real-world settings.
@@ 6.2. Auto-bidding
Auto-bidding systems have become a cornerstone of programmatic advertising (he2024hierarchical), enabling automated real-time bidding (sayedi2018real; liu2022real; cai2025rtbagent) (RTB) for ad placements while optimizing key performance indicators (KPIs) such as click-through rates (CTR), conversions, or return on ad spend (ROAS) under strict budget constraints. These systems leverage machine learning and optimization techniques to adjust bids in complex, competitive auction environments dynamically.
Reinforcement learning (RL) has gained significant traction in optimizing auto-bidding strategies (korenkevych2024offline; li2024trajectory; chen2023model). The field has evolved from early simulated bidding environments to (he2021unified)’s unified approach handling multiple constraints, followed by (jin2018real)’s multi-agent competitive settings and (mou2022sustainable)’s performance improvements over rule-based strategies. Most recently, (wen2022cooperative) deployed a cooperative multi-agent framework incorporating richer market information. Despite demonstrated successes, these RL approaches continue to face challenges, including sample inefficiency, training instability, and adaptation difficulties in dynamic auction environments.
@@ 6.3. Generative Models
Generative models aim to capture the underlying data distribution of a given dataset or model conditional probability distributions between variables. Recently, transformer-based architectures have been adapted to model complex DMP through autoregressive generative frameworks, as demonstrated in models like GAVE (gao2025generative) and GAS (li2024gas). Concurrently, diffusion models have emerged as a powerful paradigm for conditional sample generation. These models iteratively refine samples via a reverse denoising process conditioned on inputs, with representative examples including DiffBid (guo2024generative).